he coefficient of elasticity of supply of a commodity is 4. A seller supplies 40 units of this commodity at a price of birr 20 per unit. What quantity of this commodity will the seller supply when price rises by birr 24 per unit


1
Expert's answer
2022-01-05T11:13:51-0500

"E=4" "Q1=40" "p=20" "Q2=?"

"E=\frac{\delta Q}{\delta p}*\frac{p}{Q1}"

"4=\frac{\delta Q}{4}*\frac{20}{40}"

"\delta Q=32"

"32+40=72" quantity of this commodity will the seller supply